OpenStudy (anonymous):

value of cos and sin 53 in ratio

OpenStudy (anonymous):

you could basically just do this with a calculator : cos 53 =0.602 (3 significant figures) sin 53 = 0.799 ( 3 significant figures) so cos 53/sin 53 = 0.754 ( 3 sig figs)